# Technology, Institution, and Regional Growth: Evidence from Mineral Mining Industry in Industrializing Japan

### v12

@@ 5.3 Fertility Responses
The natural growth channel is testable using the birth and death statistics comprehensively measured in the vital statistics. If the population grows naturally from family planning, municipalities with mines should have had higher marriage and fertility rates. If the substitution effect were dominant, fertility rates in the mining area would be lower than those in the surrounding agrarian area (Section 4). Table 3 presents the results for the crude marriage and fertility rates from the same estimation strategy used in Table 2. Columns 1-2 and 3-4 show the results for the 1930 and 1935 samples, respectively. [^39] I compare the magnitudes for different years to discuss the potential mechanism behind the growth of the estimated magnitude of the coal mines on the local population. Note again that although several coal mines were newly opened by 1935, most already existed in 1930. Thus, the cross-sectional results from the different years can be comparable because the sample composition is similar over time. All the regressions include the fixed effects as well as the control variables.
@@
Column 1 indicates that the estimate for the crude marriage rate is negative and statistically significant. This may reflect that the influx of single male miners could be associated with lower marriage rates among the entire local population. The estimate for the crude birth rate is also negative and statistically significant. This is consistent with the lower marriage rate in the mining area and the theoretical prediction that a relatively higher wage rate for female miners before the enforcement of regulation decreases the fertility rate. To disentangle both pathways, I use the marital fertility rate as the dependent variable in column 3, which can rule out the attenuation from the greater proportion of single males in the mining area. The estimated coefficient is statistically significantly negative. The estimate ($-36.79$) is more than one standard deviation of the dependent variable, suggesting the economically meaningful effect.
@@
The results for the 1935 sample listed in columns 4–6 show a clear change after the enforcement of the female labor regulation. While column (1) shows that the estimate for the marriage rate is still statistically significantly negative, its magnitude is roughly half of the estimate for the 1930 sample. Similarly, the estimate for the crude fertility rate is roughly a quarter (column 2). Importantly, the estimate for marital fertility shows the largest magnitude reduction from the estimate in column 3. It follows that the estimates for the fertility rates are no longer statistically significant.
@@
The results show that marriage and fertility rates have reverted to their natural trends after the regulation was enforced. This aligns with the prediction described in Section 4, supporting the evidence that the declines in female relative wages had shifted women to take on the role of building a family. [^40] The optimal ratio of home labors is given by $\frac{h_{m}}{h_{f}}=\left(\frac{\kappa_{m}}{\kappa_{f}}\frac{w_{f}}{w_{m}}\right)^{\frac{1}{\rho}}$ (Online Appendix B). The labor force participation rate is only available in the 1930 Population Census. Online Appendix D.4 confirms that the mining area had lower female labor force participation than the surrounding agrarian area, with the structural shift from the agricultural to the mining sector. This is consistent with the fact that the mining area has lower female relative wages than the agrarian area (Section 2.2), supporting the labor supply mechanism behind the fertility choice. Although the data on the labor force participation rate are unavailable in the 1935 Population Census, the time-series statistics also confirm that the revision of the regulations decreased the number of female miners (Section 2.2). Under limited substitutions with the husband’s home labors and purchased services, the increase in the gender wage gap pushed the females out of the local labor market and shifted them to domestic work. This would force women to take on the role of building a family.
@@ 5.4 Mortality Changes
Table 4 presents the results for the mortality channel by gender. [^41] The result presented in this section is unchanged if I use the log-transformed rates instead of the raw rates (not reported). Columns 1–2 show the results for the 1930 census, whereas columns 3–4 show those for the 1935 sample. As shown, the estimates for the former sample are close to zero and statistically insignificant, suggesting that the mining area has a mortality rate similar to that of the surrounding area. There are a few possible interpretations of this. One possibility is that the heavy work burden in the surrounding agrarian area had balanced the mortality risk. If this channel is more relevant, the female mortality rate in the mining area could decrease after the enforcement of regulation because it does not improve the health status of the agrarian workers but works to improve the health status of the female miners. Another explanation is that the higher consumption levels offset the greater risk of miners’ mortality by improving overall health status in the mining area. If this channel is closer to reality, the enforcement should not improve mortality in the mining area because it may work to reduce the overall earnings of mining households.
Turning to look at columns 3–4, the estimate for the female mortality rate alone is weakly statistically significantly negative. The estimate is $-2.010$, roughly half of the standard deviation of female mortality in 1935. Thus, this may support the former occupational hazard channel. The institutional change from the enforcement of regulation liberated women from dangerous underground work, which may have improved the health status of females.
@@
The result for overall mortality suggests that the regulation mitigated the female miners’ occupational hazards. To illustrate a more comprehensive view of the improvements in local mortality risk, I assess to what extent occupational hazards influenced the reproductive health of female miners and whether institutional change improved such hazards. To do so, I use the early-life mortality measured in the registration-based vital statistics. I also assess the potential influence of mining pollution in the mining area.
@@
First of all, I illustrate a theoretical framework for mortality selection in utero to better interpret the quantitative evidence from the vital statistics data. The biological sorting mechanism in utero suggested by the Trivers-Willard hypothesis (TWH) (Trivers and Willard 1973) implies that there are two possible cases in which fetuses are exposed to health shocks: scarring and selection.424242 Valente (2015) tests the TWH using the case of the civil conflict in Nepal and shows that fetal exposure to the war decreased the sex ratio at births.
@@
The propositions can be summarized as follows. First, fetal health shocks may increase the risk of fetal death through both mechanisms. Second, the risk of infant mortality increases (decreases) in the scarring (selection) mechanism. Third, if both mechanisms work simultaneously, the risk of infant mortality remains unchanged. Considering these, I use the fetal death and infant mortality rates to test whether coal mines had adverse health effects and which mechanism worked in the case of mineral mining exposure.
@@
Panel A of Table 5 presents the results for the 1933 sample. Column 1 shows that the estimate for the IMR is positive and statistically significant, indicating that coal mines increased the infant mortality risk. In column 2, I run the same regression by replacing the IMR with FDR to assess the mechanism behind the higher risk of infant mortality in coal mining areas. The estimate is positive and statistically significant. This implies that the greater risk was associated with reducing fetal health endowments by mining, supporting evidence on the scarring mechanism before birth.
Next, I test whether the occupational hazard led to this scarring due to coal mining. In column 3, I limited my sample to $93$ municipalities with coal mines to analyze the correlation between the FDR and the number of female miners and accessibility to rivers. If heavy manual work during pregnancy affects fetal health, FDRs should be positively correlated with the number of female miners. To control for the scale effect and provide a placebo test, I included the number of male workers simultaneously in the same specification. The estimated coefficient for female miners is significantly positive, whereas that for male miners is rather negative. This result supports the evidence that occupational hazards for females (not males) alone increased the risk of death before birth. The estimate indicates that a one standard deviation increase in female miners ($=269.2$ miners) increases the number of fetal deaths by roughly $10$ ($0.038\times 269$) per 1000 births. This is economically meaningful because the mean difference in the FDR between the treatment and control groups was also approximately $9$ fetal deaths per 1000 births (Panel B of Online Appendix Table C.4). Column 4 shows the result for the IMR, showing that the estimated coefficient on the number of female miners is also statistically significantly positive. It further provides evidence that this reduction in fetal health endowment via occupational hazards for females is associated with higher infant mortality risks.
@@
The decline in the estimated magnitude of these health outcomes between 1933 and 1938 could partly be associated with an improving trend in health-related risk-coping strategies, such as increments in the number of medical doctors and installation of modern water supply (Online Appendix Figure D.5). However, the city and town fixed effects and accessibility variables can control for the heterogeneities in the local sanitary levels. Taken together, my results imply that the occupational hazards had increased the early-life mortality risk through the scarring mechanism in utero. However, the risks substantially decreased in the late 1930s because enforcing the revised labor regulations changed the labor patterns of females and reduced the number of female miners. Therefore, the institutional change may have attenuated the mortality selection mechanism before births throughout the 1930s.
@@
Section 2 suggests that the coal sludge generated from the coal selection process might have been a potential pollution factor. Suppose such wastewater increases the health risk of fetuses and infants. In that case, municipalities with mines closer to rivers should have significantly higher mortality rates because the wastewater had been discharged into the rivers (Chiba and Yamada 1964). However, the estimated coefficients of the river accessibility variable are close to zero and statistically insignificant in columns 3 and 4 in both panels. This result suggests that, although wastewater was occurring, as anecdotes suggest, it was not enough to harm the health conditions of mothers and infants.
Moreover, air pollution is also unlikely to increase mortality risks. If air pollution increases the early-life mortality, then the estimated coefficients on the male miner should be positive in columns 3 and 4 because larger mines should emit more coal smoke from the boilers of the steam winding machines. As shown, however, the estimates are very close to zero in all the cases. Column 5 of Panel A further tests the potential influence of air pollution using an alternative outcome, the mortality rate of children aged 1–4. If air pollution mattered, the child mortality rate would be greater in the coal mining area than in the surrounding regions because younger children are more susceptible to pollutants and are more likely to play outside than infants. The estimate is, however, close to zero and is statistically insignificant. Column 6 of Panel A uses data on $93$ municipalities with coal mines and provides further evidence: the estimated coefficients on the number of female and male miners are close to zero, suggesting that child mortality did not depend on the scale of emissions. The results for child mortality in 1938 are materially similar to those in 1933 (columns 5–6 in Panel B). Overall, air pollution does not seem to be a plausible channel to explain the greater early-life mortality in the mining area. This is consistent with the historical fact that winding machines were electrified in the interwar periods (Section 2).
@@ 5.5 Social Changes
The fertility responses and mortality changes found in Sections 5.3 and 5.4 indicate that the natural change in the mining area was initially marginally positive given the lower fertility and mortality indifferent to the surrounding area. After the enforcement of the revised regulation, the natural change may have become moderately positive because fertility reverted to a natural trend, and mortality slightly declined. Since population dynamics is the sum of natural and social growth in the short-run, social change is positively determined without exception under the clear population growth in the mining area. This means that estimating the number of immigrants is not relevant to understanding the mechanism behind the local population growth. [^45] As explained in Section 4, the systematic statistics on the number of net migrants at the municipal level were unavailable in prewar Japan. Despite this, I analyze to what extent the influx of new miners led to gender bias and influenced household size in this subsection. This is particularly useful to see how the fertility responses confirmed in Section 5.3 had influenced the local demographic characteristics in the mining area.
@@
Table 6 presents the results for the sex ratio and average household size. Columns 1–2 show the results for the 1930 sample, whereas Columns 3–4 show the results for the 1935 sample. The estimated coefficients in Columns 1 and 2 indicate that municipalities with coal mines experienced an increase in the sex ratio and a decline in the average household size. The estimated magnitudes are economically meaningful. Both estimates ($0.12$ and $0.54$) account for more than one standard deviation in the sex ratio and the average household size. This confirms that the influx of miners, many of whom were single males, led to a disproportionate sex ratio and smaller household sizes. [^46] This tendency in the mining area is in line with the phenomenon observed in the urban areas in prewar Japan. The urban areas had a relatively higher sex ratio than localities due to the internal migration of male workers (Ito 1990, p.243–247). Columns 3–4, however, show much smaller estimates for the 1935 sample: the estimates for the sex ratio and average household size are $0.07$ and $0.49$. While both estimates are still statistically significant, the large declines in magnitude are relevant to the reverting fertility after the regulation (Section 5.3). The increased fertility shall reverse the overall sex ratio because the sex ratio at birth is generally balanced at that time. [^47] See for example Ogasawara (2022) for the details about the sex ratio at birth during the prewar period. It also increases the average household size of existing households, attenuating the impacts of single immigrants.
@@
To summarize, the foregoing results provide suggestive evidence that the miners had begun to form their families in the mid-1930s in response to the labor regulations of the female miners. The total amount of miners had been relatively stable between 1930 and 1935 (Table 1). Therefore, a possible explanation for the greater marginal effect on the local population in the mid-1930s in my main result (Table 2) is an increase in the natural growth term by the local family formation and declines in the mortality risk through the institutional change in the early 1930s.
@@ 5.6 Robustness
I ensure the robustness of my results in the following ways. First, I used the cluster-robust variance-covariance matrix estimator in Online Appendix D.3. This estimator provides a systematically larger estimate for the standard error as it allows for the within-cluster dependencies. Despite this, the results are materially similar, meaning that my baseline estimates are robust to the potential influences of the local-scale spatial correlations. Second, I consider the regional heterogeneity in the endowment of male miners. The coal mines in Hokkaidō (northernmost island) were less likely to use female miners in the pits, given the nature of the labor supply of male miners. My results are materially similar if I include an indicator variable for Hokkaidō prefecture in the regressions (Online Appendix D.7). Third, I employ an alternative identification strategy. As explained in the data section, the nature of the assignments makes it difficult to employ the estimation strategy utilizing within variations of each unit. Nevertheless, I provide a case study using the DID approach for the subsample of the Tōhoku region (northeastern region). The results are reasonable to my main findings (Online Appendix D.8). Finally, I assess the validity of my results on coal mines by comparing the results for heavy metal mines. Since coal mining requires many more workers than heavy metal mining because coal is a bulky mineral, the estimated magnitudes for the heavy metal mines shall be much smaller. I confirm that the results are reasonable to such expectation (Online Appendix D.9).

### v14

@@ 5.3 Fertility Responses
The natural growth channel is testable using the birth and death statistics comprehensively measured in the vital statistics. If the population grows naturally from family planning, municipalities with mines should have had higher marriage and fertility rates. If the substitution effect were dominant, fertility rates in the mining area would be lower than those in the surrounding agrarian area (Section 3). Table 3 presents the results for the crude marriage and fertility rates from the same estimation strategy used in Table 2. Columns 1-3 and 4-6 show the results for the 1930 and 1935 samples, respectively. [^40] The magnitudes for different years are compared to discuss the potential mechanism behind the growth of the estimated magnitude of the coal mines on the local population. Note again that although several coal mines were opened by 1935, most already existed in 1930. Thus, the cross-sectional results from the different years are comparable because the sample composition is similar over time. All the regressions include the fixed effects as well as the control variables.
@@
Column 1 indicates that the estimate for the crude marriage rate is negative and statistically significant. This may reflect that the influx of single male miners could be associated with lower marriage rates among the entire local population. The estimate for the crude birth rate is also negative and statistically significant (column 2). This is consistent with the lower marriage rate in the mining area and the theoretical prediction that a relatively higher wage rate for female miners before the enforcement of regulation decreased the fertility rate. To disentangle both pathways, the marital fertility rate was used as the dependent variable in column 3 to rule out the attenuation from the greater proportion of single males in the mining area. The estimated coefficient is statistically significantly negative. The estimate ($-36.79$) is more than one standard deviation of the dependent variable, suggesting an economically meaningful effect.
@@
The results for the 1935 sample listed in columns 4–6 show a clear change after the enforcement of the female labor regulation. While column (4) shows that the estimate for the marriage rate is still statistically significantly negative, its magnitude is roughly half of the estimate for the 1930 sample. Similarly, the estimate for the crude fertility rate is roughly a quarter (column 5). Importantly, the estimate for marital fertility in column (6) shows the largest magnitude reduction from the estimate in column 3. It follows that the estimates for the fertility rates are no longer statistically significant.
@@
The results show that marriage and fertility rates have reverted to their natural trends after the regulation was enforced. This aligns with the prediction described in Section 4, supporting the evidence that the declines in female relative wages had shifted women to take on the role of family building. [^41] The optimal ratio of home labor is given by $\frac{h_{m}}{h_{f}}=\left(\frac{\kappa_{m}}{\kappa_{f}}\frac{w_{f}}{w_{m}}\right)^{\frac{1}{\rho}}$ (Online Appendix B). The labor force participation rate is only available in the 1930 Population Census. Online Appendix D.4 confirms that the mining area had lower female labor force participation than the surrounding agrarian area, with a structural shift from the agricultural to the mining sector. This is consistent with the fact that the mining area has lower female relative wages than the agrarian area (Section 2.2), supporting the labor supply mechanism behind the fertility choice. Although the data on the labor force participation rate are unavailable in the 1935 Population Census, the time-series statistics also confirm that the revision of the regulations decreased the number of female miners (Section 2.2). Under limited substitutions with the husband’s home labor and purchased services, the increase in the gender wage gap pushed the females out of the local labor market and shifted them to domestic work. This forced women to take on the role of building a family.
@@ 5.4 Mortality Changes
Table 4 presents the results for the mortality channel by gender. [^42] The result presented in this section is unchanged if the log-transformed rates is used instead of the raw rates (not reported). Columns 1–2 show the results for the 1930 census, whereas columns 3–4 show those for the 1935 sample. As shown, the estimates for the former sample are close to zero and statistically insignificant. This suggests that the mining area has a mortality rate similar to that of the surrounding area. There are a few possible interpretations of this. One possibility is that the heavy work burden in the surrounding agrarian area had balanced the mortality risk. If this channel is more relevant, the female mortality rate in the mining area could decrease after the enforcement of regulation because it does not improve the health status of the agrarian workers but improves only the health status of the female miners. Another explanation is that the higher consumption levels offset the greater risk of miners’ mortality by improving overall health status in the mining area. If this channel is closer to reality, the enforcement should not improve mortality in the mining area because it may work to reduce the overall earnings of mining households.
Turning to look at columns 3–4, the estimated coefficient for the female mortality rate alone is negative and weakly statistically significant. The estimate is $-2.010$, roughly half of the standard deviation of female mortality in 1935. Thus, this may support the former occupational hazard channel. The institutional change from the enforcement of regulation liberated women from dangerous underground work, which may have improved the overall health status of females.
@@
The result for overall mortality suggests that the regulation mitigated the female miners’ occupational hazards. To illustrate a more comprehensive view of the improvements in local mortality risk, the extent to which occupational hazards influenced the reproductive health of female miners and whether institutional change improved such hazards are assessed. This is done using the early-life mortality measured in the registration-based vital statistics. The potential influence of mining pollution in the mining area is also assessed.
@@
First of all, a theoretical framework is illustrated for mortality selection in utero to interpret the quantitative evidence from the vital statistics data. The biological sorting mechanism in utero suggested by the Trivers-Willard hypothesis (TWH) (Trivers and Willard 1973) implies that there are two possible cases in which fetuses are exposed to health shocks: scarring and selection.434343 Valente (2015) tests the TWH using the case of the civil conflict in Nepal and shows that fetal exposure to the war decreased the sex ratio at births.
@@
The propositions can be summarized as follows. First, fetal health shocks may increase the risk of fetal death through both mechanisms. Second, the risk of infant mortality increases (decreases) in the scarring (selection) mechanism. Third, if both mechanisms work simultaneously, the risk of infant mortality remains unchanged. Considering these, the fetal death and infant mortality rates are used to test whether coal mines had adverse health effects and which mechanism worked in the case of coal mining operations.
@@
Panel A of Table 5 presents the results for the 1933 sample. Column 1 shows that the estimate for the IMR is positive and statistically significant, indicating that coal mines increased the infant mortality risk. In column 2, the same regression is run by replacing the IMR with FDR to assess the mechanism behind the higher risk of infant mortality in coal mining areas. The estimate is positive and statistically significant. This implies that greater risk was associated with reducing fetal health endowments by mining, supporting evidence on the scarring mechanism before birth.
Next, we test whether the occupational hazard led to this scarring due to coal mining. In column 3, the sample is limited to $93$ municipalities with coal mines to analyze the correlation between the FDR and the number of female miners and accessibility to rivers. If heavy manual work during pregnancy affects fetal health, FDRs should be positively correlated with the number of female miners. To control for the scale effect and provide a placebo test, the number of male workers was included in the same specification. The estimated coefficient for female miners is significantly positive, whereas that for male miners is negative. This result supports the evidence that occupational hazards for females (not males) alone increased the risk of death before birth. The estimate indicates that a one standard deviation increase in female miners ($=269.2$ miners) increases the number of fetal deaths by roughly $10$ ($0.038\times 269$) per 1000 births. This is economically meaningful because the mean difference in the FDR between the treatment and control groups was also approximately $9$ fetal deaths per 1000 births (Panel B of Online Appendix Table C.4). Column 4 shows the result for the IMR, showing that the estimated coefficient on the number of female miners is also statistically significantly positive. It further provides evidence that this reduction in fetal health endowment via occupational hazards for females is associated with higher infant mortality risks.
@@
The decline in the estimated magnitude of these health outcomes between 1933 and 1938 is noteworthy. Although this could partly be associated with an improving trend in health-related risk-coping strategies such as increments in the number of medical doctors and installation of modern water supply (Online Appendix Figure D.5), the city and town fixed effects and accessibility variables can control for the heterogeneities in the local sanitary levels. Therefore, my results imply that the occupational hazards had increased the early-life mortality risk through the scarring mechanism in utero. However, the risks substantially decreased in the late 1930s because enforcing the revised labor regulations changed the labor patterns of females and reduced the number of female miners. Consequently, the institutional change may have attenuated the mortality selection mechanism before births throughout the 1930s.
@@
Section 2 suggests that the coal sludge generated from the coal selection process could have been a potential pollution factor. If such wastewater increases the health risk of fetuses and infants, municipalities with mines closer to rivers should have significantly higher mortality rates because the wastewater had been discharged into the rivers (Chiba and Yamada 1964). However, the estimated coefficients of the river accessibility variable are close to zero and statistically insignificant in columns 3 and 4 in both panels. This result suggests that, while wastewater was occurring, it was not suffincient to harm the health conditions of mothers and infants, as anecdotes suggest (Section 2.3).
Moreover, air pollution is also unlikely to increase mortality risks. If air pollution increases early-life mortality, then the estimated coefficients on the male miner should be positive in columns 3 and 4. This is because larger mines should emit more coal smoke from the boilers of the steam winding machines. As shown, however, the estimates are very close to zero in all the cases. Column 5 of Panel A further tests the potential influence of air pollution using an alternative outcome, the mortality rate of children aged 1–4. If air pollution was a contributory factor, the child mortality rate would be greater in the coal mining area than in the surrounding regions. This is because younger children are more susceptible to pollutants and are more likely to play outside than infants. The estimate is, however, close to zero and is statistically insignificant. Column 6 of Panel A uses data on $93$ municipalities with coal mines and provides further evidence. The estimated coefficients on the number of female and male miners are close to zero, suggesting that child mortality did not depend on the scale of emissions. The results for child mortality in 1938 are materially similar to those in 1933 (columns 5–6 in Panel B). Overall, air pollution does not appear to be a plausible channel to explain the greater early-life mortality in the mining area. This is consistent with the historical fact that winding machines were electrified during the interwar period (Section 2.3).
@@ 5.5 Social Changes
The fertility responses and mortality changes found in Sections 5.3 and 5.4 indicate that the natural change in the mining area was initially marginally positive given the low fertility and mortality differences with the surrounding area. After the enforcement of the revised regulation, the natural change may have become moderately positive because fertility reverted to a natural trend, and mortality slightly declined. Since short-run population dynamics is the sum of natural and social changes, social change shall be positive under the clear population growth in the mining area. Thus, estimating the number of immigrants is no longer relevant to understanding the mechanism behind the local population growth. Despite this, the extent to which the influx of new miners led to gender bias and influenced household size was analyzed in this subsection. This is useful to provide insight on how the fertility responses confirmed in Section 5.3 could be associated with the local demographic characteristics in the mining area.
@@
Table 6 presents the results for the sex ratio and average household size. Columns 1–2 show the results for the 1930 sample, whereas Columns 3–4 show the results for the 1935 sample. The estimated coefficients in Columns 1 and 2 indicate that municipalities with coal mines experienced an increase in the sex ratio and a decline in the average household size. The estimated magnitudes are economically meaningful. Both estimates ($0.12$ and $0.54$) account for more than one standard deviation in the sex ratio and the average household size, respectively. This confirms that the influx of miners, many of whom were single males, led to a disproportionate sex ratio and smaller household sizes. [^46] This tendency in the mining area is in line with the phenomenon observed in the urban areas in prewar Japan. The urban areas had a relatively higher sex ratio than rural localities due to the internal migration of male workers (Ito 1990, p.243–247). Columns 3–4, however, show much smaller estimates for the 1935 sample. The estimates for the sex ratio and average household size are $0.07$ and $0.49$. While both estimates are still statistically significant, the large declines in magnitude are relevant to the reverting fertility after the regulation (Section 5.3). The increased fertility reverses the overall sex ratio because the sex ratio at birth is generally balanced at that time. It also increases the average household size of existing households, attenuating the impacts of single immigrants.
@@
To summarize, the foregoing results provide evidence suggesting that the miners had begun to form their families in the mid-1930s in response to the labor regulations targeted at female miners. The total number of miners had been relatively stable between 1930 and mid-1930s (Table 1), suggesting that the social changes in the mining areas were stable. Consequently, a possible explanation for the greater marginal effect on the local population growth in the mid-1930s in the main result (Table 2) is that an increase in the natural growth term by the local family formation and declines in the mortality risk through the institutional change in the early 1930s.
@@ 5.6 Robustness
The robustness of the results is ensured in the following ways. First, the cluster-robust variance-covariance matrix estimator was used in Online Appendix D.3. This estimator provides a systematically larger estimate for the standard error by allowing for within-cluster dependencies. Despite this, the results are materially similar, meaning that the baseline estimates are robust to the potential influences of the local-scale spatial correlations. Second, the regional heterogeneity in the endowment of male miners is taken into account. The coal mines in Hokkaidō (northernmost island) were less likely to use female miners in the pits, given the nature of the labor supply of male miners. The results are materially similar if we include an indicator variable for Hokkaidō prefecture in the regressions (Online Appendix D.7). Third, an alternative identification strategy is employed. As explained in the data section, the nature of the assignments makes it difficult to employ the estimation strategy utilizing within variations of each unit. Nevertheless, the case study is provided using the DID approach for the subsample of the Tōhoku region (northeastern region). The results are reasonable to the main findings (Online Appendix D.8). Finally, the validity of the results on coal mines is assessed by comparing the results for heavy metal mines. Since coal mining requires many more workers than heavy metal mining because coal is a bulky mineral, the estimated magnitudes for the heavy metal mines shall be much smaller. It is confirmed that the results are reasonable to such expectation (Online Appendix D.9).